OpenStudy (anonymous):

The coordinates of C are (x,5), the coordinates of D are (6,9) and the coordinates of the midpoint, M, of Line segment CD are (-1,7). Find X.

OpenStudy (angela210793):

well the midpoint's coordinates are XM=(X1+X2)/2 and YM=(Y1+Y2)/2 can u solve this equation?